Home‎ > ‎Crew 134‎ > ‎

118-GreenHabReport

Greenhab Report
01/18/2014
Prepared by Ashley Dale

The GreenHab is operating normally

The fan was set to 60 degrees at 7:30AM
The fan is operating normally at 68 degrees at 5PM

The heater was turned down to 40 degrees F at 7:30AM
The heater was turned up to 75 degrees F at 5PM

The pump is operating normally.

The onion seedlings have sprouted. The other seedilings have note sprouted. The Mars shoot is shown some wear. 
The herb seedlings and the Mars shoot received water this morning and this afternoon. 

The hydroponic lettuce sprouts did not receive water today at 9AM. All the lettuce shoots appear healthy. 
At 5pM the hydroponic soil appeared and felt damp. The sprouts will be watered tomorrow. 
 
They remaining lettuce seedlings look very healthy. They are close to out growing the root cubes. 
The pepper seedlings have not sprouted as yet. All the seedlings in root cubes received water at 8AM. 
All the seedlings were returned to the upper hab at 5PM. 

No other plants received water today. 

Greenhab readings for 5PM today are as follows:

                    Current         Min          Max
Temperature        __ Degrees F  __ Degrees F __ Degrees F   
Humidity           __ %          __ %         __  %

Meter            1.04 KwH
Duration        24.23 Hours
Overflow            0 Liters    Removed Today at 5PM - Checked at 7:30AM and at 5PM

Regards,
End Transmission